Determine the impulse response coefficient of a low pass FIR filter of length 3 defined by a symmetrical impulse response, h[0] h[2]. Let the input to this filter be a sum of two cosine sequence of angular frequency 0.2 rad/samples and 0.4 rad/samples [Tentukan pekali sambutan dedenyut untuk suatu penapis lulus rendah (FIR) dengan panjang 3 yang ditakrifkan oleh suatu sambutan dedenyut simetri, h[0] h[2]. Biarkan masukkan kepada penapis ini sebagai hasiltambah dua urutan kosin dengan frekuensi sudut 0.2 radsampel dan 0.4 rad/sampel. (8 Marks IMarkah)
Answer #1

Answer:-

0.2 rad/s and 0.6 rad/s

thus the input is

x(t)=cos(0.2t)+cos(0.4t)

put t=nT

x(nT)=cos(0.2nT)+cos(0.4nT)

for T=1

x(n)=cos(0.2n)+cos(0.4n)

Given that the filter has symmetric impulse response with

h[0] = h[2]

The Z-transform of the given input function is

X(z)=[z(z-cos0.2)/(z^2-2zcos0.2+1)]+[z(z-cos0.4)/(z^2-2zcos0.4+1)]